Check whether the sequence (an), where


an = 1/ (n+1) + 1/(n+2) +....+1/(2n) is convergent or not




1
Expert's answer
2022-06-09T14:09:40-0400

It converges to "\ln2".


The sequence "x_n" is "H_{2n}-H_n", where "H_n" is the "n^{th}" harmonic number. It is known that "\lim\limits_{n\to+\infty}(H_n-\ln n)" is "\gamma", the Euler-Mascheroni constant. Hence

"\lim\limits_{n\to\infty}((H_{2n}-\ln2n)-(H_n-\ln n))=\gamma-\gamma=0".

"\lim\limits(H_{2n}-H_n)=\ln2n-\ln n=\ln\frac{2n}{n}=\ln2".